A fair die is rolled 10 times. What is the probability that an odd number (1, 3 or 5) will occur less than 3 times?

A. .0547
B. .1172
C. .1550
D. .7752
E. .8450
Answer #1

A. .0547
Answer #2

(1/2)^10 + 10C1*(1/2)*(1/2)^9 + 10C2*(1/2)^2*(1/2)^8 = 0.0547


Option A
Answer #3
10 C 3 + 10 C2 + 10 C 1 * 1/2^10

= (120+45+10) / 2^10

= 175/1024
= 0.17089